# A class of forward-backward regularizations of the Perona-Malik equation with variable exponent

### v1

@@ 1 Introduction
When $p>1$ is a constant, equation (1) arises in [1] as a model for image denoising, where $u$ denotes the restored image, $f$ the noisy observation, and the function values quantify the gray levels. For sufficiently small $\delta$, this equation is of forward-backward parabolic type and admits no classical strong or distributional solutions. In [2], Young measure theory was employed to study the Dirichlet initial-boundary value problem (IBVP) for equation (1) with $p=2$, where the authors proved the existence of Young measure solutions using Rothe’s method and variational techniques combined with relaxation theory [3]. Demoulini [4] further investigated the qualitative properties of these solutions, including uniqueness, the comparison principle, and long-time behavior, which were later extended from the Dirichlet to the periodic and Neumann settings in [5]. The well-posedness of the Dirichlet IBVP for equation (1) was generalized to the case $1<p\leq 2$ by adopting an approach similar to that in [4]. For the case $p=1$, equation (1) becomes singular, and Young measure solutions are considered in $L^{\infty}(0,T;BV(\Omega))$; see [6]. In fact, equation (1) with constant exponent $p>1$ serves as a forward-backward regularization of the well-known Perona-Malik equation
@@
which is parabolic if $\left|\nabla u\right|<1$ and backward-parabolic if $\left|\nabla u\right|>1$. Originally introduced for image reconstruction, the Perona-Malik equation is designed to reduce noise while preserving important image features, particularly edges and contours [7]. In practice, however, the Perona-Malik equation suffers from the drawback that it may result in staircase artifacts in the restored image. This issue inherently stems from the ill-posedness of (5), whose instabilities during the evolution manifest as the so-called staircasing phenomenon in numerical discretizations [8]; see Figure 1. From a dynamical perspective, this phenomenon is replaced by a milder, microstructured ramping in the $p$-Laplacian regularized Perona-Malik equation, characterized by alternating gradients of finite amplitude; see Figure 2.
@@
where the heat flux $\vec{\varPhi}=\nabla\varPsi$ for some nonconvex potential $\varPsi\in C^{1}(\mathbb{R}^{N})$. In this case, the monotonicity condition $(\vec{\varPhi}(\xi)-\vec{\varPhi}(\eta))\cdot(\xi-\eta)\geq 0$ is violated for some $\xi,\eta\in\mathbb{R}^{N}$. The one-dimensional version of equation (6) reads
@@
and appears independently as a model for heat and mass transfer in a stably stratified turbulent shear flow [9]. Setting $w=u_{x}$, formal differentiation of (7) with respect to $x$ leads to another type of forward-backward parabolic equation
@@
which arises in the theory of phase transitions [10]. In [11], the Neumann IBVP for (7) is shown to admit infinitely many weak $L^{2}$-solutions when $\varPhi$ is a nonmonotonic piecewise linear function satisfying $s\varPhi(s)\geq Cs^{2}$ for some $C>0$. This condition on $\varPhi$ was later relaxed in [12] by assuming higher regularity of the initial data. In [13], the existence of infinitely many weak $W^{1,\infty}$-solutions to the one-dimensional Perona-Malik equation under homogeneous Neumann boundary conditions was established. As shown in [14, 15], no global $C^{1}$ solution exists for sufficiently large $\left|f^{\prime}\right|$, which suggests that even the local existence of solutions to the IBVP for equations (6) or (7) remains a nontrivial problem. Regarding the uniqueness of forward-backward parabolic problems, the situation becomes more subtle (see [16]).
@@
From the viewpoint of the calculus of variations, the study of parabolic equations of the form (6) is closely related to the minimization of the nonconvex energy functional
$\int_{\Omega}\varPsi(\nabla u)dx.$ (9)
As far as we know, common approaches to treating functional (9) include:
  • (i)
    Convexification. The nonconvex functional can be convexified by suitable means to obtain a convex optimization problem. For instance, following the relaxation theory, the potential $\varPsi$ in (9) is replaced by its convex envelope $\varPsi^{\ast\ast}$ (see [3, 17, 18, 19, 20]).
  • (ii)
    Variational regularization. A regularizing functional, typically of higher-order or strongly convex type, is incorporated into (9) to suppress energy oscillations arising from the nonconvexity of $\varPsi$, reduce the effect of local minima, and improve stability (see [5, 1, 21, 22]). Representative examples include $\left\|\Delta u\right\|_{L^{2}(\Omega)}^{2}$ and $\left\|\nabla u\right\|_{L^{p}(\Omega)}^{p}$, the latter corresponding to equation (1) with constant $p>1$.
  • (iii)
    Generalized minimizers. Under admissible initial and boundary conditions, a family of approximation problems for the gradient flow (6) is constructed, with the expectation that their solutions, if possible, converge in some sense to a "minimizer" of functional (9), which can be captured by Young or Radon measures (see [23, 24, 25, 26]). In fact, the call for sensible generalizations of a minimizer in variational principles is analogous to the notion of a "weaker solution" to a forward-backward problem.
The so-called approximation equation appears as a suitable regularization of equation (6), often motivated by physical considerations, but it generally destroys the original variational structure. Moreover, the qualitative behavior of solutions is highly sensitive to the type of regularization employed. Here, we are mainly concerned with the latter two treatments of the nonconvex functional (9), which have been explored in various contexts (see [27, 2, 21, 28, 29, 22, 30, 31, 32]).
The pseudo-parabolic regularization of equation (7),
@@
with $\varepsilon>0$ a small parameter and $\psi^{\prime}>0$, has been proposed to account for time-delay effects [9] and is applied to modeling the formation of layers with constant temperature or salinity in the ocean [33]. By the transformation $w=u_{x}$, equation (10) is formally related to
@@
which arises in population dynamics [34]. Equation (11) serves as the pseudo-parabolic regularization of equation (8). If $\psi(u)=u$, this regularization is referred to as the Sobolev regularization [35]. The well-posedness of the Neumann IBVP for (10) was established in [33], where $\psi$ is an increasing but bounded function. It was also shown that solutions may develop spatial discontinuities (jumps) within finite time, with amplitudes that increase over time. This phenomenon results not only from the shape of $\varPhi$ but also from the boundedness of $\psi$, whereas it does not occur if $\psi$ has power-type or logarithmic growth at infinity [36, 37]. The existence of suitably defined nonnegative solutions to the Dirichlet IBVP for (11) was proved in the space of Radon measures (see [38, 39, 40, 41]), and results concerning uniqueness can be found in [42]. As $\varepsilon\to 0^{+}$, the family of solutions to the regularized problem is expected to converge to a measure-valued solution of the original IBVP for (8); see [43].
@@
where $\vec{\varPhi}$ satisfies appropriate growth conditions, were investigated in [23] (see also [44]). By the theory of Young measures (e.g., see [45, 46, 47]), it has been proved that a subsequence of the family of approximating solutions to the regularized problem (12) converges, in a suitable topology, to a Young measure solution of the ill-posed original IBVPs for (6). When the growth assumptions on $\vec{\varPhi}$ are relaxed, the Dirichlet IBVP for equation (6) was studied in [25] through the limit $\varepsilon\rightarrow 0^{+}$ of solutions to the corresponding IBVP for Sobolev regularization
@@
where the term $\Delta u_{t}$ can be interpreted as accounting for viscous relaxation effects [48]. Specifically, the first step is to address equation (13) under the Dirichlet boundary condition for any $\varepsilon>0$, and then to examine the vanishing viscosity limit $\varepsilon\rightarrow 0^{+}$ of the family $\{u^{\varepsilon}\}_{\varepsilon}$ of solutions to the approximating problems, obtaining Young measure solutions for the original problem. The asymptotic behavior of such solutions for large times is also studied in [25] using compactness arguments and $\omega$-limit set techniques. In view of the above results, a widely accepted strategy for handling forward-backward parabolic problems is to approximate them by a family of regularized problems, whose solutions are then used to extract limiting points, yielding suitably defined solutions of the original problem (see [26, 49, 50, 43]). It should be noted that, except in specific cases (see [4, 51]), the uniqueness of such solutions remains an open problem.
@@
which provides a mild variational regularization of the convex-concave energy functional underlying the Perona-Malik flow ($\delta=0$), while any $\delta>0$ is sufficient to make (14) convex-concave-convex. Figures 13 illustrate the transition from rough to smooth staircasing. In contrast to (1) with a constant exponent, the solution to the proposed regularization for a suitable choice of $p$ shows better regularity; its gradients exhibit fewer oscillations, and the maximum gradients are confined to a smaller bounded interval. Meanwhile, equation (1) can be regarded as an independent anisotropic diffusion model in image processing, where the spatially varying exponent $p$ is designed according to practical requirements. In this sense, the new model serves as an interpolation between two classical models: the Perona-Malik model and that introduced in [52]. However, its application to image processing is not the primary focus of the present paper; interested readers are referred to [53, 54, 55, 56] for further details on adaptive diffusion-based and variational models.
@@
To our knowledge, only a few papers address forward-backward parabolic problems with nonstandard growth conditions. Inspired by the idea in [25, 26], in this paper we study problem (1)–(3) using Sobolev regularization. Namely, we first investigate the existence and uniqueness of weak solutions to the homogeneous Neumann IBVP for the regularized equation
@@
where $\varepsilon\in(0,1)$. We then analyze the limit as $\varepsilon\rightarrow 0^{+}$ of the family $\{u^{\varepsilon}\}_{\varepsilon}$ of weak solutions to the above regularized problems and establish the existence of Young measure solutions to the original problem (1)–(3); see Definition 2.4. The uniqueness of such a solution is left as an open problem. We emphasize here that, although the treatment in [6] (corresponding to equation (1) with constant exponent $1<p<2$; see also [4]) involves Young measures, the spirit of that approach, which does not rely on a regularization scheme motivated by physical considerations, differs from the present one. Indeed, the approximation procedure in [6] combines the explicit time discretization for the evolution equation with variational methods used to solve the stationary problem obtained at each time step. The difficulties arising from the nonconvexity of the potential are addressed via relaxation theory (see [2] for details), whereas the feasibility of our approach is based on the observation that the discretized counterpart of the damping term $\Delta u_{t}$ provides the convexity of the stationary problem even when the corresponding potential is nonconvex (see Lemma 3.2). However, the existence of such solutions for problem (1)–(3) cannot be established following the method in [6], and, consequently, our work also serves as a theoretical complement to the studies in [6, 5, 26]. Our main result is stated as follows:
Theorem 1.1 (Existence of Young measure solutions).
@@
The rest of this paper is organized as follows. Section 2 recalls some preliminaries on variable exponent spaces and Young measures, and introduces a family of regularized problems approximating problem (1)–(3), together with the definition of Young measure solutions. Section 3 establishes the existence and uniqueness of Young measure solutions to the regularized problem. Finally, Section 4 presents the proof of Theorem 1.1.
@@ 2.1 Variable exponent spaces
There are many examples of John domains. Every ball is clearly a 1-John domain, and every bounded domain satisfying the uniform interior cone condition is a John domain. In particular, any bounded Lipschitz domain is a John domain, where a domain with $C^{0,1}$-boundary is referred to as a Lipschitz domain.
For later use, we state the following important properties of the spaces $L^{p(\cdot)}(\Omega)$ and $W^{1,p(\cdot)}(\Omega)$. Detailed discussions can be found in [57, 58, 59, 60].
Lemma 2.1.
[57] Let $p\in\mathcal{P}(\Omega)$ with $p^{-}<\infty$. If $\varrho_{L^{p(\cdot)}(\Omega)}(u)>0$ or $p^{+}<\infty$, then
@@
Lemma 2.2.
[57] Let $p\in\mathcal{P}(\Omega)$. Then $L^{p(\cdot)}(\Omega)$ and $W^{k,p(\cdot)}(\Omega)$ are Banach spaces, separable if $p$ is bounded, and reflexive and uniformly convex whenever $1<p^{-}\leq p^{+}<\infty$.
Lemma 2.3 (Hölder’s inequality).
[57] Let $p,q,s\in\mathcal{P}(\Omega)$ be such that $\frac{1}{s(y)}=\frac{1}{p(y)}+\frac{1}{q(y)}$ for a.e. $y\in\Omega$. Then, for all $u\in L^{p(\cdot)}(\Omega)$ and $v\in L^{q(\cdot)}(\Omega)$,
@@
Lemma 2.4.
[57] Let $\Omega$ be a Lipschitz domain, and let $p\in\mathcal{P}^{\log}(\Omega)$ satisfy $1\leq p^{-}\leq p^{+}<\infty$. Then $C^{\infty}(\bar{\Omega})$ is dense in $W^{k,p(\cdot)}(\Omega)$, $k\in\mathbb{N}$.
Lemma 2.5 (Poincaré inequality).
[57] Let $p\in\mathcal{P}^{\log}(\Omega)$. If $\Omega$ is a bounded $\alpha$-John domain, then
@@
By the weak $\ast$ convergence of $\{f^{m}\}_{m=1}^{\infty}$ (see [61]), there exists $N_{i}\in\mathbb{N}$ such that, for all $m>N_{i}$, ităholdsăthat
@@
Fix $N=\max_{1\leq i\leq n}N_{i}$. For any $\xi\in\varPsi$, there exists $j\in\{1,2,\cdots,n\}$ such that $\left\|\xi-\xi_{j}\right\|_{L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))}<r$. Consequently,ăforăall $m>N$, it follows from Lemma 2.3 that
@@
This completes the proof. ∎
Adapting the argument of Theorem 3.20 in [3], the following sufficient condition for weak lower semicontinuity on $W^{1,p(\cdot)}(\Omega)$ can be derived. The proof is omitted for brevity.
Lemma 2.7.

### v2

@@ 1 Introduction
When $p>1$ is a constant, equation (1) arises in [Guidotti20131416] as a model for image denoising, where $u$ denotes the restored image, $f$ the noisy observation, and the function values quantify the gray levels. For sufficiently small $\delta$, this equation is of forward-backward parabolic type and admits no classical strong or distributional solutions. In [Kinderlehrer1992001], Young measure theory was employed to study the Dirichlet initial-boundary value problem (IBVP) for equation (1) with $p=2$, where the authors proved the existence of Young measure solutions using Rothe’s method and variational techniques combined with relaxation theory [Dacorogna2008]. Demoulini [Demoulini1996376] further investigated the qualitative properties of these solutions, including uniqueness, the comparison principle, and long-time behavior, which were later extended from the Dirichlet to the periodic and Neumann settings in [GUIDOTTI20123226]. The well-posedness of the Dirichlet IBVP for equation (1) was generalized to the case $1<p\leq 2$ by adopting an approach similar to that in [Demoulini1996376]. For the case $p=1$, equation (1) becomes singular, and Young measure solutions are considered in $L^{\infty}(0,T;BV(\Omega))$; see [Yin2003659]. In fact, equation (1) with constant exponent $p>1$ serves as a forward-backward regularization of the well-known Perona-Malik equation
@@
which is parabolic if $\left|\nabla u\right|<1$ and backward-parabolic if $\left|\nabla u\right|>1$. Originally introduced for image reconstruction, the Perona-Malik equation is designed to reduce noise while preserving important image features, particularly edges and contours [Perona1990629]. In practice, however, the Perona-Malik equation suffers from the drawback that it may result in staircase artifacts in the restored image. This issue inherently stems from the ill-posedness of (5), whose instabilities during the evolution manifest as the so-called staircasing phenomenon in numerical discretizations [Guidotti2015131]; see Figure 1. From a dynamical perspective, this phenomenon is replaced by a milder, microstructured ramping in the $p$-Laplacian regularized Perona-Malik equation, characterized by alternating gradients of finite amplitude; see Figure 2.
@@
where the heat flux $\vec{\varPhi}(\xi)=\nabla\varPsi(\xi)$ for some nonconvex potential $\varPsi\in C^{1}(\mathbb{R}^{N})$. In this case, the monotonicity condition $(\vec{\varPhi}(\xi)-\vec{\varPhi}(\eta))\cdot(\xi-\eta)\geq 0$ is violated for some $\xi,\eta\in\mathbb{R}^{N}$. The one-dimensional version of equation (6) reads
@@
and appears independently as a model for heat and mass transfer in a stably stratified turbulent shear flow [Barenblatt1993341]. Setting $w=u_{x}$, formal differentiation of (7) with respect to $x$ leads to another type of forward-backward parabolic equation
@@
which arises in the theory of phase transitions [Brokate2012]. In [Hollig1983299], the Neumann IBVP for (7) is shown to admit infinitely many weak $L^{2}$-solutions when $\varPhi$ is a nonmonotonic piecewise linear function satisfying $s\varPhi(s)\geq Cs^{2}$ for some $C>0$. This condition on $\varPhi$ was later relaxed in [ZHANG2006322] by assuming higher regularity of the initial data. In [Zhang2006171], the existence of infinitely many weak $W^{1,\infty}$-solutions to the one-dimensional Perona-Malik equation under homogeneous Neumann boundary conditions was established. As shown in [GURTIN1996178, Kichenassamy19971328], no global $C^{1}$ solution exists for sufficiently large $\left|f^{\prime}\right|$, which suggests that even the local existence of solutions to the IBVP for equations (6) or (7) remains a nontrivial problem. Concerning the uniqueness of forward-backward parabolic problems, the situation becomes more subtle (see [LAIR1985311]). For more theoretical results related to the Perona-Malik equation, the reader is referred to [BOTTAZZI202139, CORLI2022474, ELLIOTT1985257, KIM20151889].
@@
A natural approach to handling ill-posed problems is to introduce suitable regularization, often motivated by physical considerations, which leads to well-posed problems [Guidotti2015131]. Nevertheless, the qualitative behavior of solutions is highly sensitive to the type of regularization employed. In recent years, various regularization schemes have been investigated in different contexts (see [BELLETTINI2008892, Giacomelli2019374, GUIDOTTI20094731, Shao20221779]). The pseudo-parabolic regularization of equation (7),
@@
with $\varepsilon>0$ a small parameter and $\psi^{\prime}>0$, has been proposed to account for time-delay effects [Barenblatt1993341] and is applied to modeling the formation of layers with constant temperature or salinity in the ocean [Barenblatt19931414]. By the transformation $w=u_{x}$, equation (9) is formally related to
@@
which arises in population dynamics [Horstmann2004545]. Equation (10) serves as the pseudo-parabolic regularization of equation (8). If $\psi(u)=u$, this regularization is referred to as the Sobolev regularization [Victor1998457]. The well-posedness of the Neumann IBVP for (9) was established in [Barenblatt19931414], where $\psi$ is an increasing but bounded function. It was also shown that solutions may develop spatial discontinuities (jumps) within finite time, with amplitudes that increase over time. This phenomenon results not only from the shape of $\varPhi$ but also from the boundedness of $\psi$, whereas it does not occur if $\psi$ has power-type or logarithmic growth at infinity [Bertsch20131719, Bertsch2016051]. The existence of suitably defined nonnegative solutions to the Dirichlet IBVP for (10) was proved in the space of Radon measures (see [BERTSCH2015217, BERTSCH2016190, BERTSCH20172037, BERTSCH201846]), and results concerning uniqueness can be found in [Smarrazzo2012085]. As $\varepsilon\to 0^{+}$, the family of solutions to the regularized problem is expected to converge to a measure-valued solution of the original IBVP for (8); see [Smarrazzo2013551].
@@
where $\vec{\varPhi}$ satisfies appropriate growth conditions, were investigated in [Slemrod1989001] (see also [Plotnikov1997550]). By the theory of Young measures (e.g., see [ball2005version, Giaquinta1998, Valadier1990]), it has been proved that a subsequence of the family of approximating solutions to the regularized problem (11) converges, in a suitable topology, to a Young measure solution of the ill-posed original IBVPs for (6). When the growth assumptions on $\vec{\varPhi}$ are relaxed, the Dirichlet IBVP for equation (6) was studied in [THANH20141403] through the limit $\varepsilon\rightarrow 0^{+}$ of solutions to the corresponding IBVP for Sobolev regularization
@@
where the term $\Delta u_{t}$ can be interpreted as accounting for viscous relaxation effects [Novick1991331]. Specifically, the first step is to address equation (12) under the Dirichlet boundary condition for any $\varepsilon>0$, and then to examine the vanishing viscosity limit $\varepsilon\rightarrow 0^{+}$ of the family $\{u^{\varepsilon}\}_{\varepsilon}$ of solutions to the approximating problems, obtaining Young measure solutions for the original problem. The asymptotic behavior of such solutions for large times is also studied in [THANH20141403] using compactness arguments and $\omega$-limit set techniques. In view of the above results, a widely accepted strategy for handling forward-backward parabolic problems is to approximate them by a family of regularized problems, whose solutions are then used to extract limiting points, yielding suitably defined solutions of the original problem (see [Evans20041599, Oliver20031380, Smarrazzo20101046, Smarrazzo2013551]). It should be noted that, except in specific cases (see [Demoulini1996376, Mascia2009887]), the uniqueness of such solutions remains an open problem.
@@
which provides a mild variational regularization of the convex-concave energy functional underlying the Perona-Malik flow ($\delta=0$), while any $\delta>0$ is sufficient to make (13) convex-concave-convex. Figures 13 illustrate the transition from rough to smooth staircasing. In contrast to (1) with a constant exponent, the solution to the proposed regularization for a suitable choice of $p$ shows better regularity; its gradients exhibit fewer oscillations, and the maximum gradients are confined to a smaller bounded interval. Meanwhile, equation (1) can be regarded as an independent anisotropic diffusion model in image processing, where the spatially varying exponent $p$ is designed according to practical requirements. In this sense, the new model serves as an interpolation between two classical models: the Perona-Malik model and that introduced in [CHEN20061383]. However, its application to image processing is not the primary focus of the present paper; interested readers are referred to [CHEN20061383, Guo2012958, Li20243374, SHAO2020103166] for further details on adaptive diffusion-based and variational models.
@@
To our knowledge, only a few papers address forward-backward parabolic problems with nonstandard growth conditions. Inspired by the idea in [Oliver20031380, THANH20141403], in this paper we study problem (1)–(3) using Sobolev regularization. Namely, we first investigate the existence and uniqueness of weak solutions to the homogeneous Neumann IBVP for the regularized equation
@@
where $\varepsilon\in(0,1)$. We then analyze the limit as $\varepsilon\rightarrow 0^{+}$ of the family $\{u^{\varepsilon}\}_{\varepsilon}$ of weak solutions to the above regularized problems and establish the existence of Young measure solutions to the original problem (1)–(3); see Definition 4. The uniqueness of such a solution is left as an open problem. We emphasize here that, although the treatment in [Yin2003659] (corresponding to equation (1) with constant exponent $1<p<2$; see also [Demoulini1996376]) involves Young measures, the spirit of that approach, which does not rely on a regularization scheme motivated by physical considerations, differs from the present one. Indeed, the approximation procedure in [Yin2003659] combines the explicit time discretization for the evolution equation with variational methods used to solve the stationary problem obtained at each time step. The difficulties arising from the nonconvexity of the potential are addressed via relaxation theory (see [Kinderlehrer1992001] for details), whereas the feasibility of our approach is based on the observation that the discretized counterpart of the damping term $\Delta u_{t}$ provides the convexity of the stationary problem even when the corresponding potential is nonconvex (see Lemma 10). However, the existence of such solutions for problem (1)–(3) cannot be established following the method in [Yin2003659], and, consequently, our work also serves as a theoretical complement to the studies in [GUIDOTTI20123226, Oliver20031380, Yin2003659]. Our main result is stated as follows:
Theorem 1 (Existence of Young measure solutions)
@@
The rest of this paper is organized as follows. Section 2 recalls some preliminaries on variable exponent spaces and Young measures, and introduces a family of regularized problems approximating problem (1)–(3), together with the definition of Young measure solutions. Section 3 establishes the existence and uniqueness of Young measure solutions to the regularized problem. Finally, Section 4 presents the proof of Theorem 1.
@@ 2.1 Variable exponent spaces
There are many examples of John domains. Every ball is clearly a 1-John domain, and every bounded domain satisfying the uniform interior cone condition is a John domain. In particular, any bounded Lipschitz domain is a John domain, where a domain with $C^{0,1}$-boundary is referred to as a Lipschitz domain; see also [Diening2011, Martio1979383].
For later use, we state the following important properties of the spaces $L^{p(\cdot)}(\Omega)$ and $W^{1,p(\cdot)}(\Omega)$. Detailed discussions can be found in [Diening2011, Kovavcik1991592, Mihuailescu20072929].
Lemma 1
[Diening2011] Let $p\in\mathcal{P}(\Omega)$ with $p^{-}<\infty$. If $\varrho_{L^{p(\cdot)}(\Omega)}(u)>0$ or $p^{+}<\infty$, then
@@
Lemma 2
[Diening2011] Let $p\in\mathcal{P}(\Omega)$. Then $L^{p(\cdot)}(\Omega)$ and $W^{k,p(\cdot)}(\Omega)$ are Banach spaces, separable if $p$ is bounded, and reflexive and uniformly convex whenever $1<p^{-}\leq p^{+}<\infty$.
Lemma 3 (Hölder’s inequality)
[Diening2011] Let $p,q,s\in\mathcal{P}(\Omega)$ be such that $\frac{1}{s(y)}=\frac{1}{p(y)}+\frac{1}{q(y)}$ for a.e. $y\in\Omega$. Then, for all $u\in L^{p(\cdot)}(\Omega)$ and $v\in L^{q(\cdot)}(\Omega)$,
@@
Lemma 4
[Diening2011] Let $\Omega$ be a Lipschitz domain, and let $p\in\mathcal{P}^{\log}(\Omega)$ satisfy $1\leq p^{-}\leq p^{+}<\infty$. Then $C^{\infty}(\bar{\Omega})$ is dense in $W^{k,p(\cdot)}(\Omega)$, $k\in\mathbb{N}$.
Lemma 5 (Poincaré inequality)
[Diening2011] Let $p\in\mathcal{P}^{\log}(\Omega)$. If $\Omega$ is a bounded $\alpha$-John domain, then
@@
By the weak $\ast$ convergence of $\{f^{m}\}_{m=1}^{\infty}$ (see [Zeidler1990]), there exists $N_{i}\in\mathbb{N}$ such that, for all $m>N_{i}$, it holds that
@@
Fix $N=\max_{1\leq i\leq n}N_{i}$. For any $\xi\in\varPsi$, there exists $j\in\{1,2,\cdots,n\}$ such that $\left\|\xi-\xi_{j}\right\|_{L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))}<r$. Consequently, for all $m>N$, it follows from Lemma 3 that
@@
This completes the proof.
Adapting the argument of Theorem 3.20 in [Dacorogna2008], the following sufficient condition for weak lower semicontinuity on $W^{1,p(\cdot)}(\Omega)$ can be derived. The proof is omitted for brevity.
Lemma 7